OpenStudy (anonymous):

find the tangent line approximation to f(x)=√(4+3x) near 0, and use it to approximate √(4.3). compute the percentage error.

OpenStudy (anonymous):

The tangent line of \(f(x)=\sqrt{4+3x}\)...Is this for calculus or pre-calc? My point is have you guys discussed the derivative?

OpenStudy (anonymous):

it is for calculus. i know about derivatives!

OpenStudy (anonymous):

Okay because I saw the work "approximation" and I though that this was for pre-calc approximate rates of change haha Anyways, we need to the find the derivative of \(f(x)\) so we have: \[f'(x)=\frac{3}{2\sqrt{4+3x}}\] So this derivative at zero is: \[f'(0)=\frac{3}{2\sqrt{4}}=\frac{3}{4}\]

OpenStudy (anonymous):

okay i understand so far!

OpenStudy (anonymous):

Now, to approximate \(\sqrt{4.3}\), In general, if you have a number \(\sqrt{x}\). The approximation can be found by looking for the square root nearest to it, a, and calculate: \[\sqrt{x}\approx \sqrt{a}+\frac{|x-a|}{2\sqrt{a}}\] So for \(\sqrt{4.3}\), notice that it is real close to the \(\sqrt{4}\) So then our x value is 4.3 and our a value is 4 so: \[\sqrt{4.3}\approx\sqrt{4}+\frac{|4.3-4|}{2\sqrt{4}}\approx2+\frac{0.3}{4}\approx2+\frac{3}{40}\approx2.075\] This is pretty darn close! The actual value is 2.0736

OpenStudy (anonymous):

i am a little confused. it said to use 3/4 to approximate √(4.3)

OpenStudy (anonymous):

I'm not sure... the only relation i see is that \(\frac{3}{40}=\frac{f'(0)}{10}\)The percentage error is defined as: \[\%_{ERROR}=\frac{|answer_{ACTUAL}-answer_{ESTIMATE}|}{answer_{ACTUAL}}\] This is: \[\%_{ERROR}=\frac{|\sqrt{4.3}-2.075|}{\sqrt{4.3}}\approx0.0654\%\]

OpenStudy (anonymous):

thanks!